# Sublinear bilipschitz equivalence and the quasiisometric classification of solvable Lie groups

### v1

@@ 5.E. Contribution to the Quasi-Isometric Classification
  1. (1)
    A quasiisometry between $G$ and $H$ cannot be ruled out with known tools.
@@
Using the Dehn function estimates, we can now summarize to which groups this applies. First, we extract the groups whose $\rho_{1}$ decompose as products from tables 12. We then divide them according to their Dehn function estimates. The result is Table 6.
$\rho_{1}(G)$ Dehn function
$\exp(n)$ $n^{2}$ $n^{3}$
$\mathbf{R}^{2}\times G_{3,3}$ $G_{5,16}^{0,\tau}$, $G_{5,17}^{\tau,0,1}$
$\mathbf{R}^{2}\times G_{3,5}^{\alpha}$ $G_{5,13}^{\alpha<1,0,1}$ $G_{5,13}^{\alpha>1,0,1}$
$\mathbf{R}\times G_{4,5}^{1,\beta}$ $G_{5,19}^{1,\beta<0}$, $G_{5,35}^{0,\beta<0}$ $G_{5,19}^{1,\beta>0}$, $G_{5,35}^{0,\beta>0}$ $G_{5,27}$, $G_{5,28}^{1}$, $G_{5,32}^{\alpha}$
$\mathbf{R}\times G_{4,8}$ $G_{5,20}^{0}$
$\mathbf{Heis}\times A_{2}$ $G_{5,25}^{1,0}$
$\mathbf{R}\times G_{4,9}^{1}$ $G_{5,30}^{1}$, $G_{5,37}$
Table 6. Indecomposable, completely solvable groups $G$ such that $\rho_{1}(G)$ is decomposable, and their Dehn functions.
The next step is to determine which of the factors that appear in these decompositions admit the assumptions of Theorem A. Since hyperbolic groups are coarse type I, Table 4 recovers the following fact:
@@
The groups $G_{3,3},G_{3,5}^{\alpha>0},G_{4,5}^{1,\beta>0},G_{4,9}^{\beta>0}$ are of coarse type I.
We conclude that if a group $G$ appearing in Table 6 has quadratic Dehn function then $\rho_{1}(G)$ admits the hypothesis of Theorem A.
Corollary 5.22.
@@
  • $\mathcal{G}^{2}_{3,3}:=\{G_{5,16}^{0,\tau},G_{5,17}^{\tau,0,1}\}$
  • $\mathcal{G}^{2}_{3,5}:=\{G_{5,13}^{\alpha>1,0,1}\}$
@@
If $G\in\mathcal{G}_{i,j}^{2},H\in\mathcal{G}_{k,l}^{2}$ are quasi-isometric, then $(i,j)=(k,l)$.
@@
The corollary could be deduced mostly from Theorem C, or entirely from Theorem 3.17. However for simplicity we give here a direct and short proof, using just Theorem A.
@@
Let $G$and $H$ be as in the statement of the Corollary. We need to rule out the possibility of an SBE between the factors of $\rho_{1}(G)$ and $\rho_{1}(H)$. Dimension considerations imply right away that $i=k$ (See [Pal22, Theorem B] whose substantial part is done by [HP13a]). Next, $G_{4,5}^{1,\beta}$ is not SBE to $G_{4,9}^{1}$: if $\beta\neq 2$, this follows from Proposition $3.8$ in [Pal20b]. If $\beta=2$, this follows from [Pal22, Theorem E]. Finally, $G_{3,3}$ and $G_{3,5}^{\alpha<1}$ are not SBE by the main theorem in [Pal20b]. ∎
@@
A basic consequence of Theorem A is that the number of direct factors of $\rho_{1}(G)$ is a SBE invariant of the group $G$ satisfying the assumptions of the theorem. Looking at Table 5 hints at groups for which the knowledge of whether or not they can be SBE to a direct product would be beneficial in the context of quasiisometric distinction. An example for such a family of groups is $G_{5,33}^{\alpha,\beta}$, specifically with $\alpha>0$ and $\beta>0$. We pose it as a question, and refer to 5.D for a short discussion on this family of groups.
Question 5.25.
Can $G_{5,33}^{\alpha>0,\beta>0}$ be SBE to a direct product of two infinite groups?
6. Groups quasiisometric to $\operatorname{Sol}_{5}$ proof of Proposition F
Eskin, Fisher and Whyte have conjectured that the class of virtually polycyclic groups should be quasiisometrically complete within finitely generated groups [EF10, Conjecture 1.2]. The first evidence for this came from the work of Shalom, who proved that any finitely generated group quasiisometric to a polycyclic group has a finite index subgroup with nonvanishing first Betti number [Sha04]. Since every polycyclic group contains a finite-index subgroup which is a uniform lattice in a simply connected solvable Lie group, the quasiisometry classification conjecture for completely solvable Lie groups [Cor18, Conjecture 19.25] would complement [EF10, Conjecture 1.2] in that it would complete the internal QI-classification of polycyclic group. (Note that when we restrict attention to the smaller class of virtually nilpotent groups, we face a similar picture, but while the quasiisometric classification of simply connected nilpotent Lie groups is still open, Gromov’s polynomial growth theorem [Gro81] can be taken as a replacement of the above conjecture of Eskin, Fisher and Whyte.)
Peng [Pen11a, Pen11b] made significant progress towards both conjectures. To state her theorems we make the following definition.
Definition 6.1.
Let $G$ be a standard solvable group. We say that $G$ is of class $(\mathcal{P})$ if the following holds:
  1. (1)
    $\mathrm{R}_{\exp}G$ is equal to the nilradical of $G$
  2. (2)
    $\mathrm{R}_{\exp}G$ is abelian
  3. (3)
    $G/\mathrm{R}_{\exp}G$ is abelian
  4. (4)
    $G$ is unimodular
Theorem 6.2 ([Pen11b, Corollary 5.3.7]).
If two groups of class $(\mathcal{P})$ and $(\mathcal{C}_{0})$ are quasiisometric, then they are isomorphic.
Theorem 6.3 ([Pen11b, Corollary 5.3.9]).
If a finitely generated group is quasiisometric to a group of class $(\mathcal{P})$, then it is virtually polycyclic.
Remark 6.4.
In [Pen11b], the statement of Theorem 6.2 is different and involves real parts of Jordan form of the adjoint action. This is because the groups are not supposed to be in $(\mathcal{C}_{0})$ a priori. It translates as follows: if two groups $G$ and $G^{\prime}$ of class $(\mathcal{P})$ are quasiisometric, then $\rho_{0}(G)$ and $\rho_{0}(G^{\prime})$ are isomorphic. With the current formulation it is easier to see that [Pen11b, Corollary 5.3.7] implies [Pen11b, Corollary 5.3.8].
The completely solvable groups of class $(\mathcal{P})$ and dimension less or equal $5$ are $G_{3,4}$, $G_{4,2}^{-2}$, $G_{4,5}^{-(1+\delta)/2,-(1-\delta)/2}$ for $10\leqslant\delta<$ 1, $G_{5,7}^{\alpha,\beta,\gamma}$ with $\alpha+\beta+\gamma=1$, $G_{5,9}^{-1-\delta,-1+\delta}$ with $10\leqslant\delta<$ 1, $G_{5,11}^{-3}$, $G_{5,15}^{-1}$, and $G_{5,33}^{-1,-1}$, which is the group $\mathrm{Sol}_{5}$.
Using Peng’s rigidity theorem and our work in Section 5, we can prove that the finitely generated groups quasiisometric to $\mathrm{Sol}_{5}$ are almost lattices in this group.
Proposition 6.5.
Let $\Gamma$ be a finitely generated group quasiisometric to $\mathrm{Sol}_{5}$. Then there is a finite-index subgroup $\Gamma_{0}$ in $\Gamma$ and a homomorphism $\Gamma_{0}\to\mathrm{Sol}_{5}$ with finite kernel and closed co-compact image.
Remark 6.6.
The group $\mathrm{Sol}_{5}$ does have lattices, see Remark 1.13.
Proof.
The proof is done by exhaustion. The group $\mathrm{Sol}_{5}$ is of class $(\mathcal{P})$, so that by Peng’s rigidity theorem 6.3 we know that $\Gamma$ is virtually polycyclic; let $G$ be a simply connected solvable group such that there exists a finite-index subgroup of $\Gamma$ that surjects with finite kernel onto a uniform lattice in $G$ [Rag72, Theorem 4.28]. We know that $G$ is quasiisometric to $\mathrm{Sol}_{5}$, so that $\dim G=5$ and $\operatorname{conedim}G=2$. Therefore $G$ is among the following list of groups:
$G_{5,19}^{0,\beta},\ G_{5,19}^{1,\beta},\ G_{5,20}^{0},\ G_{5,20}^{1},\ G_{5,27},\ G_{5,28}^{1},\ G_{5,30}^{1},\ G_{5,32}^{0},\ G_{5,32}^{\alpha},\ G_{5,33}^{\alpha,\beta},G_{5,34},\ G_{5,35},$
$G_{5,36},\ G_{5,37},\ G_{4,2}^{\alpha}\times\mathbf{R},G_{4,4}\times\mathbf{R},G_{4,5}^{\alpha,\beta}\times\mathbf{R},\ G_{4,7}\times\mathbf{R},\ G_{4,8}\times\mathbf{R},G_{4,9}^{0}\times\mathbf{R},\ G_{4,9}^{\beta}\times\mathbf{R}.$
We can rule out most of the groups in this list since they are not unimodular. The unimodular ones are
$G_{5,19}^{1,-2},\ G_{5,20}^{0},\ G_{5,33}^{-1,-1},\ G_{5,35}^{0,-2},\ G_{4,2}^{-2}\times\mathbf{R},\ G_{4,5}^{\alpha,\beta\colon\alpha+\beta=-1}\times\mathbf{R},\ G_{4,8}\times\mathbf{R}.$
By our work in Tables 4-5 all these except $G_{5,33}^{-1,-1}\simeq\mathrm{Sol}_{5}$ have exponential Dehn function; the Dehn function of $G$ is quadratic, since it is quasiisometric to $\mathrm{Sol}_{5}$ (by [Dru04, Theorem 1.1], Leuzinger-Pittet [LP04, Theorem 2.1], or the computation in Example 5.19, using Cornulier and Tessera’s [CT17, Theorem F]). So $G$ must be isomorphic to $\mathrm{Sol}_{5}$. ∎
Remark 6.7.
According to Peng [Pen11b, Corollary 5.3.11], a completely solvable group quasiisometric to $\mathrm{Sol}_{5}$ must be a semidirect product of the form $\mathbf{R}^{2}\ltimes\mathbf{R}^{3}$. This would allow to rule out some of the groups above without estimating Dehn functions, but it does not rule out $G_{4,5}^{\alpha,1-\alpha}\times\mathbf{R}$.
Proposition 6.8.
Let $G$ be a completely solvable group, quasiisometric to $\mathrm{Sol}_{5}$. Then $G$ is isomorphic to $\mathrm{Sol}_{5}$.
Proof.
The group $\mathrm{Sol}_{5}$ is amenable and unimodular, hence geometrically amenable, and geometric amenability is invariant under quasiisometry, so that $G$ must be geometrically amenable (See §11 in [Tes08] and especially Corollary 11.13 there). Moreover, since $G$ is completely solvable, it is geometrically amenable if and only if it is unimodular. So $G$ is unimodular as well, and belongs to the list of groups already considered in the proof of Proposition 6.5. The end of the proof is the same as that of Proposition 6.5. ∎
@@ Appendix A Completely solvable groups quasiisometric to symmetric spaces
Without loss of generality, we can assume that $H$ admits a left-invariant symmetric metric $g$, so that $(H,g)$ is isometric to the symmetric space $X$. Since $H$ is completely solvable, $X$ must be of non-compact type. $H$ acts simply transitively by isometries on $X$, so any larger connected Lie group $H^{\prime}$ of isometries of $X$ will contain non-trivial point stabilizers; the latter are compact, and a completely solvable group does not have non-trivial compact subgroups. So $H$ is maximal among the completely solvable groups of isometries of $X$. Now, $G$ is quasiisometric to $X$, therefore by [Cor18, Theorem 19.25] it has a continuous, proper, cocompact action by isometries on $X$. The kernel of this action is a compact, therefore trivial, subgroup of $G$ so that we may consider $G$ as a subgroup of $\operatorname{Isom}(X)$; it is a closed subgroup by properness of the action. Let $\widehat{G}$ be a maximal completely solvable subgroup of $\operatorname{Isom}(X)$ containing $G$. By the combination of [GW88, Theorem 1.11] and [GW88, Theorem 4.3], $H$ and $\widehat{G}$ are isomorphic. It remains to show that $G=\widehat{G}$. By [HP13a] we know that
@@
Appendix B A completely solvable group with non-split exponential radical and polynomial Dehn function
We give an example of a group for which the tool we developed in Proposition D and Proposition 4.2, namely a lower bound on the Dehn function using distortion in central extensions of solvable Lie groups, improves on previously known techniques, in particular on the bound given by distortion in central extensions of nilpotent groups and on Theorem 4.6.
Our example is a variation on [Cor08, Examples 4.1, 4.2], using [CT17, Example 1.5.4] as a building block. It is of dimension 13, and we do not claim this dimension is minimal for the properties we need. We first explain the logic of our construction, which is extracted from the discussion in the end of Section 4.D. We are looking for a group with quite a few properties: it must be completely solvable (i.e. in class $(\mathcal{C}_{0})$), it must not split with a nilpotent quotient (in particular it should not be generalized standard solvable, hence not in $(\mathcal{C}_{1})$), its Dehn function must be polynomial, and its distorted central extensions must be necessary for giving a lower bound on its Dehn function (so for example it must not admit a left-invariant nonpositively curved Riemannian metric). In low dimensions and for the groups of class $(\mathcal{C}_{0})$, the property of not splitting over the exponential radical is the hardest to come by. Cornulier’s construction [Cor08, Examples 4.1, 4.2] give such groups, and hints at how to obtain them in general. We vary his building blocks in order to assure that the group admits the other desired properties; the challenge of escaping the SOL obstruction is the main reason we chose [CT17, Example 1.5.4] as a building block for our example.
Let $\mathfrak{g}_{2}$ be the Lie algebra corresponding to the group $U\rtimes A$ presented in  [CT17, Example 1.5.4]. The Lie algebra $\mathfrak{u}=\operatorname{Lie}(U)$ is given by:
$\mathfrak{u}:=$ $\langle X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},X_{9},X_{12}\rangle$
with the nonzero brackets
$[X_{1},X_{2}]=X_{4},[X_{1},X_{3}]=X_{5},[X_{2},X_{3}]=X_{6},$
$[X_{1},X_{6}]=X_{9},[X_{3},X_{4}]=X_{12},[X_{2},X_{5}]=X_{9}+X_{12}.$
(The indices of the generators indeed skip 7, 8, 10 and 11: this is reminiscent to the fact that $\mathfrak{u}$ is the quotient of the free $3$-step nilpotent Lie algebra on $3$ generators $\{X_{1},X_{2},X_{3}\}$ by the ideal generated by $[X_{i},[X_{i},X_{j}]]$ for $i\neq j\in\{1,2,3\}$).
Let $\mathfrak{a}:=\operatorname{Lie}(A)$ be the abelian Lie algebra on two generators $\langle T_{1},T_{2}\rangle$, and $\mathfrak{g}_{2}:=\mathfrak{u}\rtimes\mathfrak{a}$ with the following non-zero brakcets:
$[T_{1},X_{1}]=-X_{1},[T_{1},X_{3}]=X_{3},[T_{1},X_{4}]=-X_{4},[T_{1},X_{6}]=X_{6},$
$[T_{2},X_{1}]=-X_{1},[T_{2},X_{2}]=2X_{2},[T_{2},X_{3}]=-X_{3},[T_{2},X_{4}]=X_{4},[T_{2},X_{5}]=-2X_{5},$
$[T_{2},X_{6}]=X_{6}$
Let $\mathfrak{fil}=\langle E_{1},E_{2},E_{3},E_{4}\rangle$ be the 4-dimensional filiform algebra, with non-zero brackets $[E_{1},E_{2}]=E_{3},[E_{1},E_{3}]=E_{4}$. Define $\mathfrak{g}_{3}:=\mathfrak{g}_{2}\times\mathfrak{fil}$. The centre of $\mathfrak{g}_{3}$ is the product of the centres of the factors, which is $\langle X_{9},X_{12},E_{4}\rangle$. Let $\mathfrak{z}$ be the $1$-dimensional subspace generated by the diagonal of the centre $Z:=X_{9}+X_{12}+E_{4}$.
Our example is the group $G$ whose Lie algebra is $\mathfrak{g}:=\mathfrak{g}_{3}/\mathfrak{z}$. We can write it in the basis
$\langle X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},X_{9},X_{12},T_{1},T_{2},E_{1},E_{2},E_{3}\rangle,$
with all non-zero brackets exactly as in $\mathfrak{g}_{3}$, except for $[E_{1},E_{3}]=-X_{9}-X_{12}$.
The group $G$ admits the following properties:
  • $\rho_{0}(G)=G$, i.e. $G$ is in $(\mathcal{C}_{0})$.
  • $U=\operatorname{R}_{\exp}G$ and the short exact sequence
    $1\to U\to G\to G/U\to 1$
    does not split. In particular $G$ is not in $(\mathcal{C}_{1})$.
  • $G$ is not generalized standard solvable.
  • $G$ does not have a nonabelian nilpotent retract.
  • $G$ has a polynomially bounded Dehn function.
  • $G$ admits a cubically distorted central extension.
We supply short reasoning for the above claims. Due to the high dimension of this group, we do not give the details of the computations. The reader may consult Section 5.C for the relevant techniques.
The nilradical of $\mathfrak{g}$ is $\mathfrak{u}+\mathfrak{fil}/\langle Z\rangle$, it splits with complement $\mathfrak{a}$ acting with only real eigenvalues. Therefore $G$ is in $(\mathcal{C}_{0})$.
The exponential radical is $\mathfrak{u}$. The quotient $\mathfrak{g}/\mathfrak{u}$ is $\mathbf{R}^{2}\times\mathfrak{heis}$, where in the above basis $\mathfrak{heis}=\langle E_{1},E_{2},E_{3}\rangle/\mathfrak{u}$ is the Lie algebra of the $3$-dimensional real Heisenberg group with central element $E_{3}\cdot\mathfrak{u}$. If $\mathfrak{u}$ did split, we would have $\mathfrak{g}=\mathfrak{u}\rtimes\mathfrak{m}$ with $\mathfrak{m}$ isomorphic to $\mathbf{R}^{2}\times\mathfrak{heis}$ and its action on $\mathfrak{u}$ would factor through the quotient. In particular, the central element of $\mathfrak{heis}$ in $\mathfrak{m}$, which is central in $\mathfrak{m}$, would act trivially on $\mathfrak{u}$. Therefore the centre of $\mathfrak{g}$ would intersect $\mathfrak{m}$ nontrivially. It can be verified however that the centre of $\mathfrak{g}$ is exactly $\langle X_{9},X_{12}\rangle\subset\mathfrak{u}$.
If, towards contradiction, $G$ was generalized standard solvable with nilpotent quotient $N$, then $\mathfrak{n}:=\operatorname{Lie}(N)$ would have to be a quotient of $\mathfrak{m}=\mathbf{R}^{2}\times\mathfrak{heis}$ (recall $G/\operatorname{R_{exp}(G)}$ is the largest nilpotent quotient of $G$). Therefore if $\mathfrak{n}$ is nonabelian, it must contain $\mathfrak{heis}$ and the same argument as above yields a contradiction. This moreover proves that $G$ does not retract to a nonabelian nilpotent group. If on the other hand $N$ was abelian, then by Proposition 3.9 $G$ would have to split over the exponential radical, which is not the case. So $G$ is not generalized standard solvable.
One may check that $G$ does not admit the SOL or $2$-homological obstructions, and therefore has a polynomial Dehn function [CT17, Theorem E].
It is easily observed that $G_{3}$ (the Lie group corresponding to $\mathfrak{g}_{3}$) is a central extension of $G$. The generator of the extension is $Z=X_{9}+X_{12}+E_{4}$, which is in $C^{3}\mathfrak{g}_{3}$ but not in $C^{4}\mathfrak{g}_{3}$, i.e. $c_{Z}=3$. By Proposition D, $\langle Z\rangle$ is $n^{3}$-distorted in $G_{1}:=\rho_{1}(G)$, hence by Proposition 4.2, the Dehn function of $G$ is bounded from below by $n^{3}$. A direct computation of the second cohomology group proves that we cannot improve this lower bound using distortion in other central extensions.
To the best of our knowledge, the most accurate evaluation of the Dehn function of $G$ prior to our work is given Theorem 4.6. Concretely, it is easy to check that $\rho_{1}(\mathfrak{g})=\mathfrak{u}\rtimes(\mathfrak{a}\times\mathfrak{heis})$, where the only difference in brackets from $\mathfrak{g}$ is that $[E_{1},E_{3}]=0$. In particular, the action of $\mathfrak{a}\times\mathfrak{heis}$ on the exponential radical $\mathfrak{u}$ is the same as in $\mathfrak{g}$ and so the corresponding Lie group $G_{1}=U\rtimes(A\times\operatorname{Heis})$ is generalized standard solvable in the sense of [CT17, Section 10.H]. Theorem 4.6 yields:
$n^{3}/\log^{e}(n)\preceq\delta_{G}(n)\preceq n^{4}\cdot\log^{e}(n).$
where $e\leq 8$ is twice the bound on the exponent of $\delta_{G}$.
Finally, we remark that $\operatorname{Kill}(\mathfrak{u})_{0}\neq 0$, so the upper bound of $n^{4}$ cannot be improved to $n^{3}$ using only [CT17, Theorem 10.H.1]. Moreover, $\rho_{1}(G)$ is not generalized tame as can be seen by drawing the weight diagram, and using [CT17, Proposition 4.B.5].

### v2

@@ 4. Applications to $5$-dimensional solvable Lie groups
An important quasiisometry invariant is the Dehn function. In a subsequent paper [GP] [^2] The full list of Lie groups and computations can currently be found in Section 5 of the arXiv preprint arXiv:2410.05042., using the work of Cornulier and Tessera [CT17], we compute the Dehn functions of all completely solvable groups of exponential growth up to dimension 5.
Based on these computations, we can summarize the contribution of our strategy to these groups.
First, we extract the simply connected solvable Lie groups $G$ such that $\rho_{1}(G)$ decomposes as a product, and then divide them according to their Dehn function. The result is Table 1. The names of the groups are from the classification found in [Mub63] (see also [PSWZ76]); the Lie group are named $G_{d,i}^{\alpha_{1},\ldots,\alpha_{r}}$ where $d$ is the dimension, $i$ is a positive integer, and $\alpha_{1},\ldots,\alpha_{r}$ are parameters (we keep the parameters in the same order as in [Mub63, PSWZ76], but we sometimes use different letters, in accordance with [GP])).
Image by $\rho_{1}$ Dehn function conedim
exponential quadratic cubic
$\mathbf{R}^{2}\times G_{3,3}$ $G_{5,16}^{0,\tau}$, $G_{5,17}^{\tau,0,1}$ 3
$\mathbf{R}^{2}\times G_{3,5}^{1/\alpha}$ $G_{5,13}^{\alpha<1,0,1}$ $G_{5,13}^{\alpha>1,0,1}$ 3
$\mathbf{R}\times G_{4,5}^{\gamma,1}$ ($\dagger$) $G_{5,19}^{1,\beta<0}$, $G_{5,35}^{0,\beta<0}$ $G_{5,19}^{1,\beta>0}$, $G_{5,35}^{0,\beta>0}$ $G_{5,27}$, $G_{5,28}^{1}$, $G_{5,32}^{\alpha}$ 2
$\mathbf{R}\times G_{4,8}$ $G_{5,20}^{0}$ 2
$\mathbf{Heis}\times A_{2}$ $G_{5,25}^{1,0}$ 4
$\mathbf{R}\times G_{4,9}^{1}$ $G_{5,30}^{1}$, $G_{5,37}$ 2
($\dagger$) $\gamma$ is either equal to $\beta$, or to $1$ if there is no parameter called $\beta$.
Table 1. Indecomposable, completely solvable groups $G$ such that $\rho_{1}(G)$ is decomposable, their Dehn functions, and the dimensions of their asymptotic cones.
The next step is to determine which of the factors that appear in the decompositions in the $\rho_{1}$ column admit the assumptions of Theorem C or at the very least Theorem 3.16. We do this in the following Lemma.
@@
Let $\alpha\in(0,1),\beta\in(0,1)$. The following hold true:
  1. (1)
    The groups $G_{3,5}^{\alpha},G_{4,5}^{\beta,1}$ are diagonal type Heintze groups with abelian derived subgroups.
  2. (2)
    $P_{1}=G_{3,3}$, $P_{2}=G_{4,5}^{1,1}$ and $P_{3}=G_{4,9}^{1}$ can be written as $P_{i}=A_{i}N_{i}$, $1\leqslant i\leqslant 3$, where $G_{i}=K_{i}A_{i}N_{i}$ is among the rank one simple Lie groups $G_{1}=\mathrm{SO}(3,1)$$G_{2}=\mathrm{SO}(4,1)$ and $G_{3}=\mathrm{SU}(2,1)$.
  3. (3)
    $G_{4,9}^{\beta}$ is a diagonal type Heintze group with the strong pointed sphere property.
Proof.
Let us check (1) with the help of the structure constants given in Table 2 and using Heintze’s characterisation [Hei74]. For $G=G_{3,5}^{\alpha}$ we may write that the Lie algebra is $\mathfrak{g}_{5,3}^{\alpha}=\mathbf{R}^{2}\rtimes_{\delta}\mathbf{R}$ where $\delta=\operatorname{ad}_{e_{3}}$ has strictly positive spectrum $\{1,\alpha\}$. Similarly $\mathfrak{g}_{4,5}^{\beta,1}=\mathbf{R}^{2}\rtimes_{\delta}\mathbf{R}$ where $\delta=\operatorname{ad}_{e_{4}}$ has strictly positive spectrum $\{1,\beta\}$ (with the eigenvalue $1$ having multiplicity $2$). As for (2), we see again from the structure constants that $\mathfrak{g}_{3,3}=\mathbf{R}^{2}\rtimes_{\delta}\mathbf{R}$ where $\delta=\operatorname{ad}_{e_{3}}=1$, $\mathfrak{g}_{4,5}^{1,1}=\mathbf{R}^{3}\rtimes_{\delta}\mathbf{R}$ where $\delta=1$, and $\mathfrak{g}_{4,9}^{1}=\mathfrak{heis}\rtimes_{\delta}\mathbf{R}$ where $\mathfrak{heis}$ is the $3$-dimensional Heisenberg Lie algebra and $\delta$ is its Carnot derivation (see e.g. [Pan89, §9.3]). As for (3), note that $\beta>0$ and $\beta\neq 1$ so that it is a diagonal Heintze group with derived subgroup isomorphic to $\mathbf{Heis}$, the foliation in its Gromov boundary along left cosets of $\langle e_{2}\rangle$ are preserved by the self-sublinear bilipschitz equivalences [Pal20b, Lemma 3.9], and so it has the strong pointed sphere property by using the exact same mechanism as in the proof of [Pal22, Lemma 4.1]. ∎
$A_{2}$ $e_{2},e_{1}]=e_{1}$.
$\mathbf{Heis}$ $e_{1},e_{2}]=e_{3}$.
$G_{3,3}$ $e_{3},e_{1}]=e_{1}$, $e_{3},e_{2}]=e_{2}$.
$G_{3,5}^{\alpha}$ $e_{3},e_{1}]=e_{1}$, $e_{3},e_{2}]=\alpha e_{2}$, $1\leqslant\alpha<1$, $\alpha\neq 0$
$G_{4,5}^{\alpha,\beta}$ $e_{4},e_{1}]=e_{1}$, $e_{4},e_{2}]=\alpha e_{2}$, $e_{4},e_{3}]=\beta e_{3}$, $1\leqslant\alpha\leqslant\beta\leqslant 1$, $\alpha\beta\neq 0$.
$G_{4,8}$ $e_{1},e_{2}]=e_{3}$, $e_{4},e_{1}]=e_{1}$, $e_{4},e_{2}]=-e_{2}$
$G_{4,9}^{\beta}$ $e_{1},e_{2}]=e_{3}$, $e_{4},e_{1}]=e_{1}$, $e_{4},e_{2}]=\beta e_{2}$, $e_{4},e_{3}]=(1+\beta)e_{3}$, $1<\beta\leqslant 1$.
Table 2. Structure constants for the factor groups appearing in $\rho_{1}(G)$, $G$ from Table 1 (whenever we do not write a bracket, it means it is 0).
We conclude that if a group $G$ appearing in Table 1 has quadratic Dehn function then $\rho_{1}(G)$ admits the hypothesis of Theorem C.
Corollary 4.2 (of Theorem C).
@@
  • $\mathcal{G}^{3}_{3,3}:=\{G_{5,16}^{0,\tau},G_{5,17}^{\tau,0,1}\}$
  • $\mathcal{G}^{3}_{3,5}:=\{G_{5,13}^{\alpha>1,0,1}\}$
@@
Let $c\in\{2,3\}$. If $G\in\mathcal{G}_{i,j}^{c},H\in\mathcal{G}_{k,l}^{c}$ are quasi-isometric, then $(i,j)=(k,l)$. Further, $\rho_{1}(G)$ and $\rho_{1}(H)$ are isomorphic.
@@
The index $c$ in the statement serves to record the dimension of the asymptotic cone, so that if $G\in\mathcal{G}_{i,j}^{c},H\in\mathcal{G}_{k,l}^{c^{\prime}}$ are quasiisometric, then $c=c^{\prime}$; we mention this in order to emphasize in which respect our corollary brings new information. If $(i,j)=(3,5)$, then one can further deduce that $G$ and $H$ are isomorphic, but at this stage, considering $\rho_{0}$ was actually sufficient to reach such a conclusion. All the groups in $\mathcal{G}_{3,3}^{3}$ are quasiisometric to each other (since they have the same image under $\rho_{0}$) so one should not hope to have the conclusion that $G$ and $H$ are isomorphic when $(i,j)=(3,5)$. On the other hand, it would be highly desirable to know if the conclusion of the corollary can be improved to a group isomorphism between $G$ and $H$ when $(i,j)\in\{(4,5),(4,9)\}$; in the latter case when $(i,j)=(4,9)$, this is related to Question A.2 in our Appendix, since $\mathbf{R}\times G_{4,9}^{1}$ has a Riemannian symmetric metric with a non-trivial Euclidean factor.
Corollary 4.4.
The groups $G_{5,19}^{1,\beta}$, $\beta\in(0,+\infty)$, whose Lie algebra have structure constants
$e_{1},e_{2}]=e_{3},\,[e_{5},e_{1}]=e_{1},\,[e_{5},e_{3}]=e_{3},\,[e_{5},e_{4}]=\beta e_{4},$
are pairwise non-quasiisometric.
@@
We computed that $\rho_{1}(G_{5,19}^{1,\beta})$ is isomorphic to $\mathbf{R}\times G_{4,5}^{1,\beta}$ (see Table 1) so that $\rho_{1}(G)$ records the parameter $\beta$ for $G\in\mathcal{G}_{4,5}^{2}$, hence by Corollary 4.2, $\beta$ is a quasiisometry invariant in this family. ∎
@@
Using Theorem 3.16 and part (3) in Lemma 4.1, we can prove that $\mathbf{R}\times G_{4,9}^{\beta}$ is never quasiisometric to $G_{5,30}^{1}$ or $G_{5,37}$ unless $\beta=1$. This was not possible with Theorem C only.
@@ Appendix A Completely solvable groups quasiisometric to symmetric spaces
Without loss of generality, we can assume that $H$ admits a left-invariant symmetric metric $g$, so that $(H,g)$ is isometric to the symmetric space $X$. Since $H$ is completely solvable, $X$ must be of non-compact type. $H$ acts simply transitively by isometries on $X$, so any larger connected Lie group $H^{\prime}$ of isometries of $X$ will contain non-trivial point stabilizers; the latter are compact, and a completely solvable group does not have non-trivial compact subgroups. So $H$ is maximal among the completely solvable groups of isometries of $X$. Now, $G$ is quasiisometric to $X$, therefore by [Cor18, Theorem 19.25] it has a continuous, proper, cocompact action by isometries on $X$. The kernel of this action is a compact, therefore trivial, subgroup of $G$ so that we may consider $G$ as a subgroup of $\operatorname{Isom}(X)$; it is a closed subgroup by properness of the action. Let $\widehat{G}$ be a maximal completely solvable subgroup of $\operatorname{Isom}(X)$ containing $G$. By the combination of [GW88, Theorem 1.11] and [GW88, Theorem 4.3], $H$ and $\widehat{G}$ are isomorphic. It remains to show that $G=\widehat{G}$. By [HP13] we know that